Your brother is working as a supervisor in a beverage bottling plant. One day, he let

you tour around the facility. You noticed that a beverage (mostly water) flows in a

pipe with a mass flow rate that would fill 220 cans every minute. Each can contain

355-mL of the liquid. At point 1 in the pipe, the gauge pressure is 152 kPa and the cross-sectional area is 8.00m². At point 2, 1.35 m above point 1, the cross-sectional area is 2.00m². Calculate the (A) mass flow rate; (B) volume fllow rate; (C) flow speeds at the two points; (D) gauge pressure at point 2.


1
Expert's answer
2021-02-04T19:33:45-0500

A)

"m_t=n\rho V=78.1~kg/min."

B)

"V_t=nV=78.1~l\/min."

C)

"v=\frac{V_t}{S},"

"v_1=\frac{78.1\cdot10^{-3}}{8}=9.76 ~\frac{mm}{min},"

"v_2=\frac{78.1\cdot10^{-3}}{2}=39.05 ~\frac{mm}{min}."

D)

"p_2=p_1-\rho gh=139~kPa."